OpenStudy (anonymous):

You have a 4-card deck containing a king, a queen, a jack, and a 10. You pick a card, keep it, and then pick another card. What is the probability that you pick exactly 1 face card? (Face cards are jacks, queens,and kings.) Please give an explanation.

OpenStudy (jhannybean):

@nincompoop

OpenStudy (anonymous):

what?

OpenStudy (jhannybean):

I am ressurecting him to assist you with your question~ Rise my minion!

OpenStudy (anonymous):

oh

OpenStudy (jhannybean):

Tbh I am horrible at stats,or rather, probability related questions.

OpenStudy (anonymous):

i just really need the explanation. i have the answer its just that i am confused on how the question is worded

OpenStudy (anonymous):

thanks for trying though

OpenStudy (nincompoop):

@waterineyes please help

OpenStudy (nincompoop):

I thought I'd learn the mathematical way from one of the best LOL

OpenStudy (anonymous):

Don't know really but let me give a try. If you pick 10 at first, then probability of face cards will be 1/3.. And if you pick one face card on first try, then probability of other face card will be 1/2. And if you pick face card on first try, then 10 at other then, no probability.. :) Don't know actually, and I don't want to misguide the asker..

OpenStudy (nincompoop):

LMAO! I was thinking with four cards, there's 3 out of 4 chances; but since we picked an unknown card, and left with 3 with the following possibilities: 3 face cards, which will give us 1 out 3 chances of getting a face;2 face cards and a 10 will give us 1 out of 3 or 2 out of 3. Now, I am stuck in putting this in mathematical equation or composition.

OpenStudy (nincompoop):

@KingGeorge I have a feeling that the new graduate will not upset us :D

OpenStudy (kinggeorge):

I would naturally think about this more in the way that waterineyes thought about it. Basically, you have two ways to do it. Case 1: You choose the 10 first. Then anything else you draw afterwards will be a face card. The probability of this happening is 1/4 (chance to draw the 10). Case 2: You choose something else first, and then the 10. This has a (3/4)*(1/3)=3/12 chance of happening. So the total probability is (3/12)+(1/4)=1/2 There is another method that I like as well.

OpenStudy (kinggeorge):

The other method would be to look at all the possible two card hands you draw. There are \(\binom{4}{2}=6\) ways to choose 2 cards. How many of these have the 10? Well, we have to have the 10, so we have three remaining cards left to choose one card from. So \(\binom{3}{1}=3\). Hence, the probability is 3/6=1/2.

OpenStudy (kinggeorge):

Did at least one of these methods make sense?

OpenStudy (nincompoop):

the second answer you gave made more sense to me. it cuts through the minute details.

OpenStudy (kinggeorge):

It certainly is a bit more elegant. But I think the first one is the more "obvious" way to do it, and perhaps a bit easier to understand.

OpenStudy (nincompoop):

agreeable. the second one is more systematically, which also a bit easier for me to understand and translate to higher magnitude or complexity.

OpenStudy (kinggeorge):

Agreed. It's much easier to generalize the second one to more difficult problems.